OpenStudy (anonymous):

2. Write the equation of the line that goes through the points (-1, 3) and (4, -7) in general form. Clearly show and explain each step @jazy ?

OpenStudy (anonymous):

can you help me ?

OpenStudy (anonymous):

\[m = \frac{ y2-y1 }{ x2-x1 }\] \[y - y1 = m(x-x1)\]

OpenStudy (anonymous):

1st find the slope and use point Slope form

OpenStudy (anonymous):

i still dont understand

OpenStudy (anonymous):

do you know the values for y in (-1, 3) and (4, -7)

OpenStudy (anonymous):

3, -7?

OpenStudy (anonymous):

Step 1) Find slope. \[m = \frac{ y2 - y1 }{ x2 - x1 } = \frac{ -7 - 3 }{ 4 - (-1) } = \frac{ -10 }{ 5 } = -2\] Step 2) Point-Slope form. Slope is 2 and you have a given point, lets use (-1, 3). \[y - y1 = m(x - x1)\]\[y - 3 = -2(x + 1)\]Now you want it in slope-intercept form. Distribute.\[y - 3 = -2x - 2\]Solve for y: Add 3 to both sides.\[y = -2x + 1\]

OpenStudy (anonymous):

Standard form: Ax + By = C Add 2x to both sides. Standard form: 2x + y = 1

OpenStudy (anonymous):

Not sure which you needed. (:

OpenStudy (anonymous):

thx jazy you make everything easier to understand

OpenStudy (anonymous):

Glad to help. (: